OpenStudy (anonymous):

if sinhy=4sinhx-3/4+3sinhx find dy/dx

OpenStudy (anonymous):

Do you know the derivative of sinhy??

OpenStudy (anonymous):

\[\frac{d}{dx}\sin(hx) = \cos(hx)\]

OpenStudy (anonymous):

i dont understand explain

OpenStudy (abb0t):

Take the derivatives of the hyperbolic trig functions. @waterineyes has given you some great information that will help you :)

OpenStudy (abb0t):

Except it's \(sinh(x) = cosh(x)\) :)